OpenStudy (anonymous):

If f(x) = −2x − 4 and g(x) = x + 4, find ()(−8)

OpenStudy (abb0t):

find g(-8) or f(-8)

OpenStudy (precal):

find ()(-8)? that notation is invalid

OpenStudy (abb0t):

Whether it is f(-8) or g(-8). What it is telling you to do is to plug in -8 for x i your function. For instance, if it said f(12345) when f(x) = 2x then you would have f(12345) = 2(12345)

OpenStudy (precal):

but it could also be a composition? so the asker needs to recheck their instructions

OpenStudy (anonymous):

If f(x) = −2x − 4 and g(x) = x + 4, find (f/9)(−8)

OpenStudy (precal):

that is f divided by g correct?

OpenStudy (anonymous):

yes :)

OpenStudy (precal):